If demand function is given as the following: 

Qz = 230 -2.75 Pz + 0.5 I + 1.2 Pm + 0.6A 

Where Qz is quantity of Good z sold, Pz is price of Good z per unit, I is per capita income, Pm is price of competitor and A is the amount of advertising spent. 

Current values:  Pz= RM 55 I= RM 9000    Pm= RM 50    A =RM 12,000

Please Determine the total revenue maximizing price and quantity


1
Expert's answer
2021-05-04T07:24:14-0400
"SOLUTION"

given demand function:

"Q_Z=230-2.75P_Z+0.5I+1.2P_M+0.6A"

And

"Current\ values: Pz= RM 55\ I= RM 9000 \ Pm= RM 50 \ A =RM 12,000"

"Q_Z=230-(2.75\times55)+(0.5\times9000)+(1.25\times 50)+(0.6\times12000)"

"Q_Z=230-151.25+4500+60+7200"

"Q_Z=11,838.75"


"TR=P\times Q"

TR="(11838.75 )p"

MR=P

P="\frac{TR}{Q}"

P="\frac{11838.75}{Q}"